What is the first line of the equation for:
-3x+2y=-6
-2x+y=6
The answer is (-18,-30) I just don't know how to write out the problem.

Answer by rfer(16322)

-3x+2y=-6
-2x+y=6
change to
y=2x+6
-----------
-3x+2*(2x+6)=-6
-3x+4x+12=-6
x=-6-12
x=-18
--------------
-3*-18+2y=-6
54+2y=-6
2y=-6-54
2y=-60
y=-60/2
y=-30
-------
(-18,-30)